OpenStudy (anonymous):

f(x) = 2 + x^(2/3) find the critical values, intervals where f(x) is increasing, intervals where f(x) is decreasing, list any max or min points

OpenStudy (anonymous):

\[f(x) = 2+ x ^{2/3)}\]

OpenStudy (anonymous):

Thank you for that!

OpenStudy (anonymous):

know of any website that plots ?

OpenStudy (anonymous):

www.wolframalpha.com

OpenStudy (anonymous):

i have that but it doesnt show how to find the rest

OpenStudy (anonymous):

sorry about asking website for plot

OpenStudy (anonymous):

first find derivative \[\frac{2}{3 x^{1/3}}\]

OpenStudy (anonymous):

at x=0 this is undefined, so this is our criticle value \[\frac{2}{3 x^{1/3}}\]

OpenStudy (anonymous):

The whole thing is never zero, so no other criticle value

OpenStudy (anonymous):

check left and right of x=0

OpenStudy (anonymous):

how to do that in this denominator

OpenStudy (anonymous):

how to do what?

OpenStudy (anonymous):

check for x=0

OpenStudy (anonymous):

Criticle point is where either derivative function is zero or undefined let's check for function=0 \[\frac{2}{3 (x)^{1/3}}==0\] function is never zero but at x=0 function is undefined

OpenStudy (anonymous):

you wrote check left and right of x = 0 where do i check this

OpenStudy (anonymous):

i am just trying to figure that out

OpenStudy (anonymous):

Hold on, I am trying to plot

OpenStudy (anonymous):

This is your function(look at blue line, ignore the red line) http://www.wolframalpha.com/input/?i=x^(2%2F3)%20%2B%202&t=crmtb01

OpenStudy (anonymous):

how is the minimum 2 when the graph extends below 2? on my graphing calculator it comes out completely different

OpenStudy (anonymous):

graph is above to 2 2 to above -inf to 2 2 to inf domain

OpenStudy (anonymous):

I see, wolfram alpha is graphing it wrong, trust your graphing calculator

OpenStudy (anonymous):

There should not be a graph left of 0

OpenStudy (anonymous):

appreciate that

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

try just graphing x^(2/3)

OpenStudy (anonymous):

wolfram alpha is seriously flawed http://www.wolframalpha.com/input/?i=cube+root+of+-8 http://tinyurl.com/3rffdsw

OpenStudy (anonymous):

it spills out good derivatives though

OpenStudy (anonymous):

Criticle point is where either derivative function is zero or undefined let's check for function=0 23(x)1/3==0 function is never zero but at x=0 function is undefined

OpenStudy (anonymous):

didnt understand the critical value part still

OpenStudy (anonymous):

i know u are trying to find where the equation is undefined how do i write the answer

OpenStudy (anonymous):

criticle value is 0 because that's where function is undefined

OpenStudy (anonymous):

cool will start a new thread shortly with another graph

OpenStudy (anonymous):

Function in red, Derivative in Blue